OpenStudy (anonymous):

Ok last question on my calc 3 homework and I can not seem to get the correct answer. Can anyone help me please?

OpenStudy (anonymous):

\[\int\limits_{0}^{2\pi}\int\limits_{0}^{\pi/6}\int\limits_{\sqrt{3}\sec \phi}^{4}3\rho^3\sin \phi d \phi d \rho d \theta\]

OpenStudy (anonymous):

whenever i solve it i get 48.433 but apparently it is incorrect

OpenStudy (anonymous):

sorry should be \[d \rho d \phi d\]

OpenStudy (amistre64):

it might be easier to use abc instead of those greek things

OpenStudy (amistre64):

the greeks makes it look like these are cylindrical coords

OpenStudy (anonymous):

ok so \[\int\limits_{0}^{2\pi}\int\limits_{0}^{\pi/6}\int\limits_{\sqrt(3)\sec(b)}^{4}3a^2\sin(b)dadbdc\]

OpenStudy (amistre64):

err, sphericals ;)

OpenStudy (anonymous):

haha ok well that is what this homework is covering spherical and cylindrical

OpenStudy (amistre64):

so lets start with the innards\ \[\int_{\sqrt(3)\sec(b)}^{4}3a^2\sin(b)da\]

OpenStudy (amistre64):

also, as a clean up, we can pull out the constants and associate them with their appropriate integrands

OpenStudy (amistre64):

\[3\int_{0}^{2\pi}\left(\int_{0}^{\pi/6}\sin(b)\left(\int_{\sqrt(3)\sec(b)}^{4}a^2da\right)db\right)dc\]

OpenStudy (anonymous):

alright looks good. the innermost one will actually cancle out with the 3 that you brought to the outside

OpenStudy (anonymous):

so from there i got \[\int\limits_{0}^{2\pi}\int\limits_{0}^{\pi/6}\sin(b)(64-(\sqrt(3))^3\sec^3(b)dbdc\]

OpenStudy (amistre64):

\[3 \int_{0}^{2\pi}\left(\int_{0}^{\pi/6}\sin(b)\left(\int_{\sqrt(3)\sec(b)}^{4}a^2da\right)db\right)dc\] \[3 \int_{0}^{2\pi}\left(\int_{0}^{\pi/6}\sin(b)\left(\frac134^3-\frac13\sqrt{3}sec(b)\right)db\right)dc\] \[\int_{0}^{2\pi}\left(\int_{0}^{\pi/6}\sin(b)\left(64-3\sqrt{3}sec^3(b)\right)db\right)dc\] \[\int_{0}^{2\pi}\left(\int_{0}^{\pi/6}64sin(b)-3\sqrt{3}tan(b)sec^2(b)~db\right)dc\] \[\int_{0}^{2\pi}\left(-64\frac{\sqrt3}{2} - \frac32 \sqrt{3}~\frac43+64+\frac{3\sqrt{3}}2\right)dc\] \[-\frac{8\sqrt3}{2} - 2\sqrt{3}+64+\frac{3\sqrt{3}}2~\int_{0}^{2\pi}~dc\] something along those lines if i aint messed it up

OpenStudy (anonymous):

ok um so that whole thing * 2pi?

OpenStudy (amistre64):

i see i ran over a step :) \[\int_{0}^{2\pi}\left(\int_{0}^{\pi/6}64sin(b)-3\sqrt{3}tan(b)sec^2(b)~db\right)dc\] \[\int_{0}^{2\pi}\left(-64cos(30^o)-\frac32\sqrt{3}~sec^2(30^o)-(-64cos(0)-\frac32\sqrt{3}~sec^2(0))\right)dc\] but yes

OpenStudy (anonymous):

weird is saying it is wrong. every answer i have tried has been wrong so i guess i will email him about it. maybe the key is wrong

OpenStudy (anonymous):

even wolframs triple integral calculator says it should be something i have tried already

OpenStudy (amistre64):

so what does the key say?

OpenStudy (amistre64):

was the integral given to you? or did you try to define it with given information

OpenStudy (anonymous):

it was given and idk what it says. its online and i only can only check if they are right or wrong. i emailed him earlier about it and he said "it looks like you have a sign switched"

OpenStudy (amistre64):

abt 48.433 is what i keep getting from what youve posted. i know some programs are picky if you dont input the answer in a specific format ..

OpenStudy (amistre64):

are you able to take a screen shot and post it as a file attachement?

OpenStudy (anonymous):

no thats not it because it converts it everthing to decimal numbers the grades the decimal and it only takes 5 places so that should be right

OpenStudy (amistre64):

does it want an approximation? or the exact stuff with srts and pis?

OpenStudy (anonymous):

either way is fine but i have tried both ways

OpenStudy (amistre64):

\[(128-65\sqrt{3})*\pi\]

OpenStudy (anonymous):

nope :/ its ok lol thank you for your help though

OpenStudy (anonymous):

im pretty sure there is just an error with the key

OpenStudy (amistre64):

GOOD LUCK ;)

OpenStudy (amistre64):

opps, caps got locked